OpenStudy (anonymous):

i need the solution set for this radical equation:: -sqrt{6x+4}=4

OpenStudy (anonymous):

no solution exists. do you know why? because the sqrt function is always non-negative.

OpenStudy (anonymous):

thats a tick mark so you could find the problem everything in the problem is positive

OpenStudy (anonymous):

oh, i see. then square both sides: 6x + 4 = 16 subtract 4 from both sides: 6x = 12 divide both sides by 6: x = 2

OpenStudy (anonymous):

and also this one: sqrt{2x}+5=13

OpenStudy (anonymous):

Checking against original equation... -4 - 4. Solution set = { }.

OpenStudy (anonymous):

she said the "-" wasn't a negative sign, it was a "tick." so the solution is x = 2

OpenStudy (anonymous):

sqrt{6x+4}=4 6x+4=16 6x=12 x=2

OpenStudy (anonymous):

If the negative sign doesn't belong there and is removed, then the solution set to the revised equation is indeed {2}.